OpenStudy (mendicant_bias):

Alright, parametrizing some basic surfaces, posted below, working on it now. Simple problems.

OpenStudy (mendicant_bias):

http://i.imgur.com/IxxHpKS.png

OpenStudy (mendicant_bias):

Guessing z = r^2, r is less than or equal to 2? Then, for the actual position vector parametrized,

OpenStudy (mendicant_bias):

\[R(r,\theta)=(r \cos(\theta), \ r \sin(\theta),\ r^2), \ \ \ 0 \leq \theta \leq 2\pi, \ \ \ 0 \leq r \leq 2\]